# Do AI Models Perform Human-like Abstract Reasoning Across Modalities?

### v3

@@ 5 Conclusions
The contributions of this work are threefold. (1) We demonstrated the effects of task representation (textual or visual), reasoning effort, and Python tool use on the ConceptARC benchmark for abstract reasoning, finding that in textual modalities with medium reasoning effort, the best AI models match or surpass humans in output accuracy. (2) We evaluated not only accuracy, but also the rules that AI models generated to describe their solutions, and found that while they were able to capture intended abstractions in about half the cases in textual settings, in other cases their rules relied on more superficial features or patterns that are less generalizable. These results suggest that relying on accuracy alone to evaluate abstract reasoning capabilities, as was done in the ARC-Prize challenge, may overestimate the generality of these capabilities. (3) We showed that state-of-the-art multimodal reasoning models still lack human-like visual reasoning abilities, performing dramatically worse in the visual than in the textual modality. However, these models were substantially better at generating correct rules than they were at applying them, which points to directions for improving visual reasoning in such systems.
Improving the abstraction capabilities of AI models is an essential direction for future research. Recognizing and using human-like abstract concepts is a crucial step for AI systems to become more generalizable and trustworthy in their reasoning, and also to successfully communicate with humans about their reasoning processes.
@@ 6 Limitations
The work we reported here has a number of limitations:
  • The results reported here involved only the ConceptARC dataset. It is possible that tasks in the original ARC test sets are more resistant to the kinds of rule “shortcuts” seen in our results; however, to our knowledge, there has been no prior research on this topic.
  • We cannot be certain that the natural-language rules generated by the AI models we evaluated are faithful representations of the actual reasoning the models do to solve a task, though in general the output grids generated seem to align with the rules. Further study is needed to quantify this alignment.
  • Due to resource limitations, we did not experiment with the “high-effort” reasoning setting for o3 or larger reasoning-token budgets for Claude and Gemini—these settings could very well produce significantly more correct-intended rules and higher accuracies.
  • Our classification of human- and machine-generated rules was done manually, and involved some subjectivity; we do not know of any objective or algorithmic means to usefully classify these natural-language rules into our various categories. However, to mitigate individual subjectivity, our team discussed and came to consensus on all potentially ambiguous classifications.
  • Due to resource limitations, we used pass@1 accuracies for both humans and machines, which differs from the pass@2 and pass@3 accuracies reported in other ARC evaluations.
  • We used the same prompt as in the ARC-Prize evaluation of o3 (Chollet, 2024) for the textual setting, and a slightly modified version for the visual setting. It may be that other prompts would elicit better performance for these systems.
  • The data we obtained for human-generated rules was not complete—no rules were collected for incorrect outputs, and even for the correct outputs, in some cases the human-generated rules were not classifiable due to reasons described earlier in the paper.
Acknowledgments
@@
Sandia National Laboratories is a multimission laboratory managed and operated by National Technology and Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International, Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA-0003525. This work was conducted as part of the BANYAN Institute, funded by Sandia National Laboratories’ lab-directed research and development program. This work was also supported by the Templeton World Charity Foundation, Inc. (funder DOI 501100011730) under the grant https://doi.org/10.54224/20650. We thank Kaleda K. Denton for helpful comments on the manuscript.
@@ Appendix B Visual Prompt
The left side of the first image shows 3 grids, where each grid square is colored with one of 10 possible colors: black, blue, red, green, yellow, gray, magenta, orange, cyan or brown. The right side of the first image also contains 3 grids, each of which is a transformed version of the corresponding grid on the left. There is a single rule that describes these 3 transformations. No Tools Variant Determine the rule, and then apply that rule to the grid in the second image. Do not generate any Python code or use any external tools to solve this task. Tools Variant Determine the rule, and then apply that rule to the grid in the second image. Use python if needed. You may describe the final grid through natural language using the indices of the different colors, for example: which would be a 6x5 grid with a horizontal blue line in row 2 and a horizontal yellow line in row 4. The rest of the grid is black. [Uncaptioned image] Image 1: Training examples [Uncaptioned image] Image 2: Test grid — Return only this minified JSON object: "rule":"<Transformation rule>","grid":"<final grid>" No markdown, no extra keys, no code fences. —
@@ Appendix C Prompts for Non-Reasoning models
Appendix D Data for Rule Evaluation Plots
Table 2: Data used to create Figure 2. For o3, Claude, Gemini, and human-generated rules, each cell reports the percentage of tasks in a rule classification (Correct-Intended, Correct-Unintended, Incorrect), partitioned by the modality (Textual vs. Visual) and by the correctness of the output grid (Correct Grid vs. Incorrect Grid). Model percentages are computed over 480 total tasks. Human percentages are computed over approximately 4,175 total tests. Rules were not collected for incorrect grids in the original experiment, and so all human responses with incorrect grids are listed here as Not Classified; these percentages are estimates based on the 73% grid accuracy reported by the original experimenters. The final row of statistics for humans show the rule classification breakdown excluding not-classified rules.
Model (Output Correctness) Textual Visual
Correct- Intended Correct- Unintended Incorrect Correct- Intended Correct- Unintended Incorrect
o3 (Correct Grid) 54.8 15.6 5.2 20.2 5.8 3.1
o3 (Incorrect Grid) 2.3 12.7 9.4 18.8 13.8 38.3
Claude Sonnet 4 (Correct Grid) 44.4 4.4 6.3 4.0 0.4 2.5
Claude Sonnet 4 (Incorrect Grid) 13.5 9.2 22.3 9.8 2.5 80.8
Gemini 2.5 Pro (Correct Grid) 42.5 10.6 7.3 4.6 0.2 1.0
Gemini 2.5 Pro (Incorrect Grid) 1.7 6.3 31.7 19.2 4.4 70.6
Correct- Intended Correct- Unintended Incorrect Not Classified
Human (Correct Grid) 53 3 3 14
Human (Incorrect Grid) 27
Human (Excl. Not-Classified) 90.0 5.1 4.9
Table 3: Data used to create Figure 3. For all o3 settings, each cell reports the percentage of tasks in a rule classification (Correct-Intended, Correct-Unintended, Incorrect), partitioned by the modality (Textual vs. Visual) and by the correctness of the output grid (Correct Grid vs. Incorrect Grid). All percentages are computed over 480 total tasks.
o3 Setting (Output Correctness) Textual Visual
Correct- Intended Correct- Unintended Incorrect Correct- Intended Correct- Unintended Incorrect
Low effort (Correct Grid) 49.4 13.1 5.8 5.3 0.2 1.0
Low effort (Incorrect Grid) 4.2 7.5 20.0 25.5 4.2 63.8
Medium effort (Correct Grid) 52.7 17.3 7.1 3.1 0.6 1.9
Medium effort (Incorrect Grid) 1.9 11.0 10.0 27.3 5.6 61.5
Low effort + tools (Correct Grid) 45.8 16.5 5.6 14.4 2.1 1.7
Low effort + tools (Incorrect Grid) 2.1 11.3 18.8 23.3 7.9 50.6
Medium effort + tools (Correct Grid) 54.8 15.6 5.2 20.2 5.8 3.1
Medium effort + tools (Incorrect Grid) 2.3 12.7 9.4 18.8 13.8 38.3
@@ Appendix E Output Accuracy for Non-Reasoning Models
As shown in Table 4 the accuracies of the non-reasoning models were dramatically lower than those of the reasoning models (Table 1). For GPT-4o, in almost all cases in both modalities, the model generated an output grid that was incorrect. For Llama 4 Scout and Qwen 2.5 VL 72B, the same was true in the textual modality; however, for Qwen, in almost all cases in the visual modality, the model was not able to generate an answer at all and did not return the requested JSON format. This was true to a lesser extent for Llama 4 Scout. It is a topic for future research to determine why these models had difficulty generating answers in any valid format.
@@
Llama 4 Scout 6.7 / 0.0 N/A
Qwen 2.5 VL 72B 9.2 / 0.0 N/A
Appendix F Concept Performance Overview
ConceptARC (Moskvichev et al., 2023) is organized around 16 basic spatial and semantic concepts. Each concept group consists of 10 tasks that focus around that concept in different ways, testing the derived knowledge on three separate test input grids. Table 5 and Table 6 give the per-concept-group accuracies (each out of 30 grids) of the reasoning models we evaluated (using medium reasoning effort and Python tools), as well as human accuracies on these concept groups from Moskvichev et al. (2023). Humans were tested using visual images of demonstration and test grids, but human accuracy is repeated in both tables for easy comparison.
F.1 Concept Performance Comparison for Textual Modality
Table 5: Concept performance (Textual): Per-concept accuracy (%) on Concept-ARC for medium effort + tools. Best value per concept in bold.
@@ F.1 Concept Performance Comparison for Textual Modality
F.2 Concept Performance Comparison for Visual Modality
Table 6: Concept performance (Visual): Per-concept accuracy (%) on Concept-ARC for medium effort + tools. Best value per concept in bold.
@@ F.2 Concept Performance Comparison for Visual Modality
F.3 Concept Difficulty Evaluation
@@
While, we do not discover a significant correlation between concept difficulty in visual or textual modality, or with human participants, we identify some overarching trends in concept difficulty. We show a full concept performance comparison in the tables Table 5 and Table 6, but specifically point out the performance differences over the concepts “Count” and “CleanUp.” We generally compare against the highest performing setting per concept for each performance. Count tasks frequently involve the production of simple, singular output rows or columns, denoting the count of specific characteristics (e.g shapes, colors, corners). Correspondingly, output grids are often small and easy to generate. In the visual modality, this is the performance closest to humans for both o3 (-7.7%) and Gemini (-44%), and in the textual modality, this concept also results in the biggest positive difference(o3:+32.3%; Gemini:+25.7%; Claude:+15.7%). In contrast, tasks in the CleanUp concept group require the removal of several colors, shapes, or isolated pixels, as well as full reproduction of the remaining input grid. In this concept group, even o3 using medium effort + tools is significantly outperformed by human participants in the visual setting (-65.7%). Similarly, answers to CleanUp tasks constitute the largest negative performance gap in the textual modality (-46.3%). The gap between the other models is even larger. This is a strong indicator that, regardless of modality, models struggle significantly with producing complex output grids.
@@
Figure 5: We show two example demonstrations from the concept with the highest and lowest gap between Human and Model performance, CleanUp and Count. Further, we show concept-wise output grid accuracy across three reasoning models in a medium with tools setting (note that we compare against the strongest setting in subsection F.3 instead).
Appendix G Correct-Intended Coverage
@@
Table 7: Correct-intended task coverage: Number of tasks covered correctly by category and modality, with coverage rates listed as a percentage of the 480 total ConceptARC tasks. Here, a task is considered "covered" if the model in question produced a correct-intended rule in any of its solutions for that task in the given modality. The "AnyModel" rows show task coverage aggregated across all three reasoning models, and the entry for humans shows the coverage of tasks for which at least one human subject produced a correct-intended rule.
@@
Humans Overall 475 98.96
o3 Textual 410 85.42
@@
Claude Textual 343 71.46
Claude Visual 80 16.67
Gemini Textual 293 61.04
Gemini Visual 136 28.33
AnyModel Textual 451 93.96
@@
G.1 Correct-Intended Coverage Implications
Table 7 clearly shows that, while models (in textual modality) all have a decent coverage, pooling their answers only leads to a moderate increase as compared to the best performing single model (+8%). While the overall coverage is notably lower in visual modality, the increase when pooling the three models again is comparable to textual (+8%). As we do not have individual human performance data, we unfortunately cannot compute similar statistics for pooling single human performances. However, these results again point out stronger abstractive reasoning abilities in a human panel, which only failed to derive the correct abstract transformation in 5 test examples.
Appendix H Error-Type Overview
Figure 6: Overview of different error types for o3 in different experimental settings. For each setting, we display textual modality results on the left bar and visual modality on the right bar. The most common error type is a simple mismatch error in which the output grid and the ground truth grid are not identical, including incorrect grid dimensions and single-pixel mismatches. We also encountered some parsing errors, which most often originated from incorrect formatting (see Appendix I). Another parsing error originated from uneven row lengths, which made it impossible to render the candidate grids in the needed rectangle format.
Appendix I Output Grid Accuracies Reassessed For Incorrect Grid Formats
To compute the accuracies reported in Table 4 and Table 1, we followed the ARC-Prize evaluation method ARC-Prize (2024): we counted an output grid as correct only if it perfectly matched the ground-truth output grid and was in the format requested in the prompt (see Appendix Aand Appendix B). However, upon exhaustive examination of the output grids generated by different models, we found that, in some cases, models generated these answer grids in different formats than that requested in the prompt; these answers were assessed as incorrect. The incorrect output grid formats included surrounding grid rows with brackets, using commas or slashes as row separators, and several other variations.
@@ Appendix I Output Grid Accuracies Reassessed For Incorrect Grid Formats
Figure 7 gives a plot corresponding to Figure 2 but with the revised accuracies. Comparing this to Figure 2, we do not see any substantial changes in the fractions of correct-intended, correct-unintended, and incorrect rules associated with each bar.
@@
Figure 7: Re-assessed rule evaluations. Results of rule evaluations, similar to that shown in Figure 2, but here with re-assessed accuracies.

### v4

@@ 6 Conclusions
The contributions of this work are threefold. (1) We demonstrated the effects of task representation (textual or visual), reasoning effort, and Python tool use on the ConceptARC benchmark for abstract reasoning, finding that in textual modalities with medium reasoning effort, the best AI models match or surpass humans in output accuracy. (2) We evaluated not only accuracy, but also the rules that AI models generated to describe their solutions, finding that while models captured the intended abstractions in about half of the cases in textual settings, in other instances their rules relied on more superficial features or patterns that are less generalizable. These results suggest that relying on accuracy alone to evaluate abstract reasoning capabilities, as was done in the ARC-Prize challenge, can substantially overestimate the generality of these capabilities. (3) We showed that state-of-the-art multimodal reasoning models still lack human-like visual reasoning abilities, performing dramatically worse in the visual modality than in the textual modality. However, in visual settings, these models were substantially better at generating correct rules than they were at applying them, pointing to potential directions for improving visual reasoning in such systems. Improving the abstraction capabilities of AI models is an essential direction for future research. Recognizing and using human-like abstract concepts is a crucial step toward making AI systems more generalizable and trustworthy in their reasoning, as well as enabling them to effectively communicate their reasoning processes to humans.
@@ 7 Limitations
The work we report here has a number of limitations. Our study involves only the ConceptARC dataset. It is possible that tasks in the original ARC test sets are more resistant to the kinds of rule “shortcuts” seen in our results; however, to our knowledge, there has been no prior research on this topic. Due to resource limitations, we did not experiment with the “high-effort” reasoning setting for o3 or larger reasoning-token budgets for Claude and Gemini. These settings could very well produce significantly more correct-intended rules and higher accuracies. In addition, our manual classification of human- and machine-generated rules involved some subjectivity; we do not know of any objective or algorithmic means to usefully classify these natural-language rules into our various categories. However, to mitigate individual subjectivity, our team discussed and came to consensus on all potentially ambiguous classifications. Also due to resource limitations, we used pass@1 accuracies for both humans and machines, which differs from the pass@2 and pass@3 accuracies reported in other ARC evaluations. Additionally, we used the same prompt as in the ARC-Prize evaluation of o3 (Chollet, 2024) for the textual setting, and a slightly modified version for the visual setting. Other prompts may elicit better performance for these systems. The data we obtained for human-generated rules were incomplete. No rules were collected for incorrect outputs, and even among correct outputs, some human-generated rules were not classifiable for reasons described earlier in the paper.
Impact Statement
Assessing the general abilities of state-of-the-art AI models to perform abstract reasoning is crucial for systems that will be deployed in many real-world settings. To date, while state-of-the-art AI models may outperform humans on abstract reasoning benchmarks, they may remain brittle on more open-ended real-world tasks.
The ARC domain has been particularly prominent as a test of human-like abstraction, and the high accuracy of models such as o3, Claude Sonnet 4, and Gemini 2.5 Pro on ARC has resulted in claims that AI models have made huge breakthroughs in human-like abstraction abilities. Our work demonstrates in detail how AI models can, in many cases, reason in ways that differ substantially from human reasoning on a version of this widely-used benchmark. We show that evaluations based only on accuracy can substantially overstate progress on these systems’ general abilities (or, in the case of visual inputs, understate such progress). Our work shows that when human-like reasoning is a desired outcome, it is essential to go beyond accuracy and examine the actual strategies used to solve tasks. In addition, our work indicates that for models to incorporate human-like priors, additional training or novel architectures specifically incorporating those priors might be required.
Given the potential value of our detailed data and results for follow-up research, we will publish a web page for this paper containing all data and code upon publication.
8 Acknowledgements
@@
Sandia National Laboratories is a multimission laboratory managed and operated by National Technology and Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International, Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA-0003525. This work was conducted as part of the BANYAN Institute, funded by Sandia National Laboratories’ lab-directed research and development program. This work was also supported by the Templeton World Charity Foundation, Inc. (funder DOI 501100011730) under the grant https://doi.org/10.54224/20650.
@@ Appendix B Visual Prompt
The left side of the first image shows 3 grids, where each grid square is colored with one of 10 possible colors: black, blue, red, green, yellow, gray, magenta, orange, cyan or brown. The right side of the first image also contains 3 grids, each of which is a transformed version of the corresponding grid on the left. There is a single rule that describes these 3 transformations. No Tools Variant Determine the rule, and then apply that rule to the grid in the second image. Do not generate any Python code or use any external tools to solve this task. Tools Variant Determine the rule, and then apply that rule to the grid in the second image. Use python if needed. You may describe the final grid through natural language using the indices of the different colors, for example: which would be a 6x5 grid with a horizontal blue line in row 2 and a horizontal yellow line in row 4. The rest of the grid is black. [Uncaptioned image] Image 1: Training examples [Uncaptioned image] Image 2: Test grid — Return only this minified JSON object: “rule”:“⟨Transformation rule⟩”,“grid”:“⟨final grid⟩” No markdown, no extra keys, no code fences. —
@@ Appendix C Prompts for Non-Reasoning Models
Appendix D Examples of Correct-Unintended Rules
In the following example, o3, using medium effort and tools, performs shallow inference for a task from the Horizontal vs. Vertical concept group. The model does not recognize the relation between the orientation of the colored shape components and the blue row, but rather focuses on whether a blue (“8”) pixel appears in the grid. In this case, the correct-intended rule works for the given test case, but does not work for other test variants.
The next example features o3, using medium effort and tools, on a task from the Complete Shape concept group. The model does not recognize the relationship between the colored output shape and the gray prototype and instead overfits to the demonstrations, producing a correct-unintended rule based on shallow features.
In the next example, Claude Sonnet 4 uses a density heuristic to approximate the most overlapped figure on a task from the Top vs. Bottom 3D group. While this approach works for some of the test examples, it fails to capture the notion of the bottommost shape in a 3D stack and does not generalize to closely related variations.
@@ Appendix E Output Accuracy for Non-Reasoning Models
As shown in Table 3 the accuracies of the non-reasoning models were dramatically lower than those of the reasoning models (Table 1). For GPT-4o, in almost all cases in both modalities, the model generated an output grid that was incorrect. For Llama 4 Scout and Qwen 2.5 VL 72B, the same was true in the textual modality; however, for Qwen, in almost all cases in the visual modality, the model was not able to generate an answer at all and did not return the requested JSON format. This was true to a lesser extent for Llama 4 Scout. Determining why these models had difficulty generating answers in a valid format is a topic for future research.
@@
Llama 4 Scout 6.7 / 0.0 -
Qwen 2.5 VL 72B 9.2 / 0.0 -
Appendix F Table 1 Output Accuracy Without Order5
We discovered that in the ConceptARC corpus (Moskvichev et al., 2023), one of the 160 tasks (task 5 of the“Order” concept group) includes a training demonstration containing a misplaced grid cell. We reanalyzed the data in Table 1 with Order5 removed and obtained the results shown in Table 4. The resulting changes are negligible and do not meaningfully affect our conclusions.
Table 4: This table reports the same data as Table 1, with Order5 removed from the analysis.
Model low medium low +T medium +T
o3 t 68.6 77.6 68.3 75.9
v 6.7 5.7 18.2 29.4
o4-mini t 52.4 70.9 57.7 77.8
v 3.8 8.2 6.7 25.2
Claude t N/A 60.6 N/A 55.3
v N/A 5.2 N/A 6.9
Gemini t N/A 66.5 N/A 60.4
v N/A 4.2 N/A 5.9
Appendix G Error-Type Overview
Figure 4: Overview of different error types for o3 in different experimental settings. For each setting, the left bar shows textual modality results and the right bar shows visual modality results. The most common error type is a simple mismatch error, in which the output grid and the ground truth grid are not identical, including cases of incorrect grid dimensions and single-cell mismatches. We also encountered parsing errors, most often originating from incorrect formatting (see Appendix K)as well as from grids with rows of differing lengths.
Appendix H Concept Performance Overview
ConceptARC (Moskvichev et al., 2023) is organized around 16 basic spatial and semantic concepts. Each concept group consists of 10 tasks that focus on the concept in different ways, with each task containing three distinct test grids. Table 5 and Table 6 give the per-concept-group accuracies (each out of 30 grids) of the reasoning models we evaluated (using medium reasoning effort and Python tools), as well as human accuracies on these concept groups from Moskvichev et al. (2023). Human trials were administered solely using visual images of the demonstration and test grids, but human accuracy is reported in both tables for ease of comparison.
H.1 Concept Performance Comparison for Textual Modality
Table 5: Concept performance (Textual): Per-concept accuracy (%) on Concept-ARC using medium effort + tools. Highest accuracy per concept is shown in bold.
@@ H.1 Concept Performance Comparison for Textual Modality
H.2 Concept Performance Comparison for Visual Modality
Table 6: Concept performance (Visual): Per-concept accuracy (%) on Concept-ARC using medium effort + tools. Highest accuracy per concept is shown in bold.
@@ H.2 Concept Performance Comparison for Visual Modality
H.3 Concept Difficulty Evaluation
@@
Although we find no significant correlation in concept difficulty across modalities (visual vs. textual) or with human participants, we do identify several overarching trends. Full per-concept performance comparisons are shown in Table 5 and Table 6; in particular, note the performance differences for the concepts “Count” and “CleanUp.”Tasks in the group “Count” frequently involve the production of simple, singular output rows or columns, denoting the count of specific characteristics (e.g shapes, colors, corners). Correspondingly, output grids are often small and easy to generate. In the visual modality, this is the performance closest to humans for both o3 (-7.7%) [^7] Note that for this and the following comparisons we consider the strongest model settings, which is not necessarily medium with tools and Gemini (-44%), and in the textual modality, this concept also results in the biggest positive difference (o3: +32.3%; Gemini: +25.7%; Claude: +15.7%). In contrast, tasks in the CleanUp concept group require the removal of several colors, shapes, or isolated pixels, as well as full reproduction of the remaining input grid. In this concept group, even o3 using medium effort + tools is significantly outperformed by human participants in the visual setting (-65.7%). Similarly, answers to CleanUp tasks constitute the largest negative performance gap in the textual modality (-46.3%). The gap between the other models is even larger. This is a strong indicator that, regardless of modality, models struggle significantly with producing complex output grids.
@@
Figure 5: Top row: Two example demonstrations from the concepts with the highest and lowest gaps between human and model performance, Count and CleanUp. Bottom row: Concept-wise output grid accuracy across three reasoning models in a medium with tools setting (note that we compare against the strongest settings in subsection H.3).
Appendix I Data for Rule Evaluation Plots
Table 7: Data used to create Figure 2. For o3, Claude, Gemini, and human-generated rules, each cell reports the percentage of tests in a rule classification (Correct-Intended, Correct-Unintended, Incorrect, or Not-Classified), partitioned by the correctness of the output grid (Correct Grid vs. Incorrect Grid), for both modalities. Model percentages are computed over a total of 480 tasks. Human percentages are computed over approximately 4,175 total tests. A rule may be labeled Not-Classified because the rule could not be clearly classified by our team, or because no rule could be collected for that particular test. Rules were not collected for incorrect grids in the original human experiment, and so all human responses with incorrect grids are reported here as Not-Classified; these statistics are estimates based on the 73% grid accuracy reported by the original experimenters. The final row of statistics for humans show the rule classification breakdown when excluding Not-Classified rules.
Model (Modality) Correct Grid Incorrect Grid
Correct- Intended Correct- Unint. Incorrect Not- Classif. Correct- Intended Correct- Unint. Incorrect Not- Classif.
o3 (Textual) 55.0 15.8 4.8 0.0 2.3 12.7 8.8 0.6
o3 (Visual) 20.4 5.6 3.1 0.0 19.6 12.9 32.5 5.8
Claude Sonnet 4 (Textual) 44.2 5.2 5.0 0.6 13.3 9.4 16.3 6.1
Claude Sonnet 4 (Visual) 4.0 0.4 2.1 0.4 10.4 2.5 66.5 13.8
Gemini 2.5 Pro (Textual) 43.8 11.9 4.8 0.0 2.3 10.0 25.6 1.6
Gemini 2.5 Pro (Visual) 4.6 0.2 1.0 0.0 19.0 5.4 67.1 2.7
Human (All Data) 53.7 2.7 3.0 13.6 27.0
Human (Excl. Not-Classified) 90.3 4.6 5.1
Table 8: Data used to create Figure 3. For all o3 settings, each cell reports the percentage of tasks in a rule classification (Correct-Intended, Correct-Unintended, Incorrect, Not-Classified), partitioned by the correctness of the output grid (Correct Grid vs. Incorrect Grid), for both modalities. A model rule may be labeled Not-Classified because the rule could not be confidently evaluated by our team, or because the model did not output a rule. All percentages are computed over 480 total tasks.
o3 Setting (Modality) Correct Grid Incorrect Grid
Correct- Intended Correct- Unint. Incorrect Not- Classif. Correct- Intended Correct- Unint. Incorrect Not- Classif.
Low effort (Textual) 49.4 13.1 5.8 0.0 4.2 7.5 18.8 1.3
Low effort (Visual) 5.4 0.2 1.0 0.0 26.3 4.4 56.7 6.0
Medium effort (Textual) 52.7 17.3 7.1 0.0 1.9 11.0 9.8 0.2
Medium effort (Visual) 3.1 0.6 1.9 0.0 27.3 5.2 54.2 7.7
Low effort + tools (Textual) 45.8 16.5 5.6 0.0 2.1 11.3 15.2 3.5
Low effort + tools (Visual) 14.4 2.1 1.7 0.0 23.3 7.9 35.8 14.8
Medium effort + tools (Textual) 55.0 15.8 4.8 0.0 2.3 12.7 8.8 0.6
Medium effort + tools (Visual) 20.4 5.6 3.1 0.0 19.6 12.9 32.5 5.8
Appendix J Correct-Intended Coverage
@@ Appendix I Data for Rule Evaluation Plots
Table 9: Correct-intended task coverage: Number of tasks covered correctly by category and modality, with coverage rates listed as a percentage of the 480 total ConceptARC tasks. Here, a task is considered “covered” if the model in question produced a correct-intended rule in any of its solutions for that task in the given modality. The “AnyModel” rows show task coverage aggregated across all three reasoning models, and the entry for humans shows the coverage of tasks for which at least one human subject produced a correct-intended rule.
@@
Humans Overall 476 99.17
o3 Textual 412 85.83
@@
Claude Textual 346 72.08
Claude Visual 83 17.29
Gemini Textual 326 67.92
Gemini Visual 140 29.27
AnyModel Textual 454 94.58
@@
J.1 Correct-Intended Coverage Implications
Table 9 shows that, while models (in textual modality) all have a decent coverage, pooling their answers only leads to a moderate increase as compared to the best performing single model (+9%). While the overall coverage is notably lower in visual modality, the increase when pooling the three models again is comparable to textual (+8%). As we do not have individual human performance data, we unfortunately cannot compute similar statistics for pooling single human performances. However, these results again point out stronger abstractive reasoning abilities in a human panel, which only failed to derive the correct abstract transformation in 4 test examples (AboveBelow2 Test 3, CompleteShape4 Test 1, HorizontalVertical5 Test 1, Order9 Test 3). It is worth noting that our human rule dataset did not contain rules for grids with incorrect outputs, meaning that the score of 476/480 is a lower bound, as some humans who produced incorrect grids may nevertheless have inferred the correct-intended rules.
Appendix K Output Grid Accuracies Reassessed For Incorrect Grid Formats
To compute the accuracies reported in Table 3 and Table 1, we followed the ARC-Prize evaluation method (ARC-Prize, 2024): we counted an output grid as correct only if it perfectly matched the ground-truth output grid and was in the format requested in the prompt (see Appendix A and Appendix B). However, upon exhaustive examination of the output grids generated by different models, we found that, in some cases, models generated these answer grids in different formats than that requested in the prompt; these answers were assessed as incorrect. The incorrect output grid formats included surrounding grid rows with brackets, using commas or slashes as row separators, and several other variations.
@@
Appendix L Distribution of Correct-Unintended Rules Across Concepts
Upon discovering the usage of correct but unintended rules, we were interested in analyzing the distribution of these among different concepts. In particular, models might be systematically employing unintended rules on tasks they lack meaningful priors for. Under textual modalities, when models arrived at a “correct” rule, they produced correct-unintended results in about 29.82% those cases with a standard deviation of 16.44%. The concepts with the highest share of correct-unintended rules were TopBottom3D (70.59%), CleanUp (52.27%) and HorizontalVertical (45.12%). For the visual modality, the average usage of correct-unintended rules was 27.0%, with a standard deviation of 13.92%. Again, the concepts with the highest share are TopBottom3D (62.50%) and CleanUp (40%), but also SameDifferent (47.83%). HorizontalVertical had a reduced share of unintended rules, with only 35.42%, ranking at fourth-most.
We generally refer to the share of unintended abstractions of correct rules here (intended and unintended), rather than of all rules, in order to account for the difference in rule correctness between modalities. As Table 5 and Table 6 show, TopBottom3D is one of the most difficult concepts for models when measured by accuracy (second-lowest in textual, third-lowest in visual), so it is not surprising that models largely rely on unintended rules when solving corresponding tasks. Analyzing the rules generated for tasks in this concept group more closely, few of them actually addressed 3-dimensional arrangement, but instead relied on shallower features, such as density or bounding-box interceptions.
While CleanUp produced the lowest overall accuracy in textual modality, it achieved the fourth-highest in visual modality, so it is intuitive to find increased shortcut-usage using text representations. The high share of unintended rules in vision-based inputs is somewhat surprising, but is likely due to difficulties with recognizing global patterns. This is an issue with both modalities, as models tend to employ local patterns, or case-by-case logic, which were overfit to the training demonstrations. In particular, models frequently recognize simple line-based patterns, such as alternating horizontal or vertical lines, but struggle with recognizing individual objects,since they seem to lack a strong prior for “objectness.”
In the other concept groups, and on various tasks in general, we found that the models’ generated rules involve recurring heuristics, including the employment of bounding boxes, four/eight-neighbor connectivity, as well as finding paths to the grid edge or object boundary by stepping through adjacent (4- or 8-connected) cells. These unintended abstractions seemed to be part of a general-purpose tool-box, which models employed for various purposes and not specific to single concepts.